# An object-centric core metamodel for IoT-enhanced event logs

### v1

@@
An object-centric core metamodel for IoT-enhanced event logs
@@ 1 Introduction
The development of Internet-of-Things (IoT) technologies has led to a widespread deployment of IoT devices by organizations to support their business processes (BPs). This trend is observed across many industries, for example, in manufacturing, where sensors and actuators pave the way for process automation (Malburg et al. 2020); in healthcare, where sensors continuously monitor the vital signs of patients to assist clinical decision-making and enforce best practices in patient handling (Fernandez-Llatas et al. 2015); in smart spaces, where sensors and actuators support the routines of elderly people (Leotta et al. 2015).
@@
As discussed by Janiesch et al. (2020), the emergence of these so-called IoT-enhanced BPs calls for the development of dedicated business process management (BPM) techniques. These should, for example, coordinate process execution with IoT devices, respond to issues detected by sensors, and provide deeper insights related to the influence of the physical environment on process execution. To achieve these goals, new IoT-enhanced process mining (PM) techniques should be capable of analyzing IoT data together with process events.
@@
One major hurdle preventing the creation of IoT-enhanced PM techniques is the need to integrate IoT data and process event data (Bertrand et al. 2022; Mangler et al. 2024). This difficulty is mainly due to the nature of IoT sensor data that is generally at a low granularity level. This in turn means that the data has to be abstracted or filtered to extract relevant information. Moreover, IoT sensor data is noisy and voluminous. This is exacerbated by the fact that IoT devices usually collect data independently of the processes. As a result, relevant IoT sensor data has to be correlated with process data to ensure a meaningful alignment between IoT observations and process events.
@@
To address this integration challenge, several models describing IoT-enhanced event log formats have recently been proposed (e.g., Mangler et al. (2023); Bertrand et al. (2023); Franceschetti et al. (2023c)). Although they share some similarities, these formats have been designed with different requirements. This is mainly due to 1) each model focusing on specific types of use cases, i.e., different types of processes and IoT devices, which generate different types of data, and 2) each model being optimized for a different data analysis task, with specific input data requirements. For example, DataStream (Mangler et al. 2023) assumes a certain degree of process knowledge, typical of manufacturing production processes, and is intended for long-term data storage for auditing. In contrast, CAIRO (Franceschetti et al. 2023c) is designed for mostly manual and more knowledge-intensive processes. Its main analysis goal is process monitoring in the absence of information system that tracks processes. In addition to this, they are grounded on different event log standards, i.e., DataStream is an extension of the eXtensible Event Stream (XES) standard (Günther and Verbeek 2014), while NICE (Bertrand et al. 2023) follows the object-centric paradigm. This discrepancy between existing IoT-enhanced log formats hinders interoperability and data reuse in research.
@@ 2 Running example
To illustrate the need for a specialized format for IoT-enhanced event logs, this section outlines a running real-life manufacturing use case, first introduced in Bertrand et al. (2024). Figure 1 shows a BPMN diagram of this process.
This use case presents the production process of a company active in the preparation of chemical products. Their production process can be summarized in four main steps: 1) Preparing raw material and loading the tank; 2) Mixing the raw material in the tank; 3) Circulating the product through filters to remove impurities; and 4) Bottling and packing the finished product.
@@
Sensors placed around the process monitor various physical parameters affected by the process, such as the tank pressure and the flow of the product through the filters. The execution of each activity leaves a specific pattern in these IoT data, hence events and activity executions can be detected by analyzing the IoT data. One such event is a peak in the flow of the product, which can indicate an anomaly in the filtering phase.
It would be useful to detect peaks in the product flow so that process engineers are able to address anomalies early and to prevent the flow from fluctuating too much, which can cause degradation of the quality of the end product. However, this requires analyzing the IoT data from the flow along with process data indicating which activity is currently being executed, a task that is not straightforward. The IoT data are at a low granularity level and need to be abstracted to gain business meaning before they can be combined with process event data (Mangler et al. 2024).
@@ 3.1 Process mining
Process mining is a method that leverages event data to analyze and enhance business processes (van der Aalst 2016). It acts as a bridge between the fields of data science and process science (van der Aalst 2022b, 2016). While data science typically does not prioritize processes and process science focuses mainly on process modeling rather than extracting insights from event data, process mining serves as a crucial tool that seamlessly combines the strengths of both disciplines (van der Aalst 2016).
@@
    Process Discovery focuses on the automatic creation of process models derived from an event log (van der Aalst 2022a).
@@
    Conformance checking, on the other hand, aims to identify deviations between real-world process events recorded in event logs and a predefined process model (Carmona et al. 2018).
@@
    Performance analysis addresses challenges such as delays in case completion or missed deadlines. By integrating event data with process models using techniques like token-based replay and alignment methods, it uncovers inefficiencies and bottlenecks (Leemans et al. 2018; van der Aalst 2022b).
@@
    Comparative process mining relies on the analysis of multiple event logs, which may differ based on attributes such as location, time period, or resource allocation. This approach allows for the identification of critical similarities and differences between processes (van der Aalst 2022b; van der Aalst et al. 2015).
@@
    In contrast to these retrospective techniques, predictive process mining combines predictive analytics with process mining methods to forecast future events, outcomes, and remaining time in support of proactive decision-making (Schulte et al. 2017; van der Aalst 2022b; Di Francescomarino and Ghidini 2022).
@@
    Finally, action-oriented process mining bridges the gap between analysis and implementation by translating insights into actionable steps to enhance operational efficiency, e.g. using robotic process automation (Park and van der Aalst 2022; van der Aalst 2022b).
@@ 3.1.1 Process Mining in IoT
IoT devices generate vast amounts of data reflecting the physical context in which processes operate, offering a rich resource for process mining. Using continuous streams of data from interconnected devices, process mining can move beyond traditional applications to provide context-sensitive, real-time insights into processes, opening new possibilities for advanced techniques (Leotta et al. 2015). For instance, process mining is increasingly employed in IoT-based systems to detect misbehavior and potential attacks (Hemmer et al. 2020), as well as to diagnose discrepancies between the designed IoT system and its actual usage. This facilitates anomaly detection and ensures that systems operate as intended (Yamaguchi and Ahmadon 2022).
The time series data produced by IoT devices often includes detailed, low-level information about processes and their executions. This enables the discovery of processes at various levels of abstraction, enhancing event log-driven process discovery by embedding IoT concepts directly into process models (Elkodssi et al. 2023). Moreover, IoT data provides valuable contextual information about the physical environments in which processes occur. Incorporating this context into process mining allows for a more in-depth understanding of the factors influencing process execution, leading to more informed decision-making (Bertrand et al. 2024).
However, integrating IoT data into process mining introduces significant challenges. First, there is a substantial gap between the granular data in sensor logs and the higher-level actions needed for process mining, further complicated by the uncertainty and noise inherent in IoT data, which may lead to ambiguity (Franceschetti et al. 2023b). Second, modeling human behavior is complex due to its unstructured and flexible nature, requiring adaptations such as fuzzy mining or declarative approaches to account for variability and probabilistic patterns. Finally, automatic log segmentation is particularly challenging as sensor logs lack predefined traces and must account for varied routines, temporal constraints, and multi-user interactions, necessitating innovative methods for accurately defining and differentiating human habits (Leotta et al. 2015).
@@ 3.2.1 XES
The current standard event log model is the eXtensible Event Stream (XES) format, an XML-based model that mainly consists of the notions of event, case, and log (Günther and Verbeek 2014). It proposes standard attribute types to contextualize the events, e.g. the resource executing an activity, the cost of an activity, etc. A standard activity lifecycle is defined together with XES, based on which the status of an activity can be mapped with events relating to this activity. XES also allows the definition of new data attribute types through extensions.
@@ 3.2.2 OCEL/OCED
Recently, the gain in maturity of the PM field has increased the urge to create alternative models to the original XES standard. Multiple propositions that relax some assumptions of XES and allow for more flexibility in event data storage have been presented, e.g., in Popova et al. (2015); Ghahfarokhi et al. (2021); Berti et al. (2024); Fahland et al. (2024). In this context, the object-centric event data (OCED) initiative (Fahland et al. 2024) led by the IEEE task force in process mining (TFPM) aims at updating the XES format used by the community with a more flexible standard for event data. The main innovation of OCED is the introduction of the concept of object, which generalizes the notion of case by allowing one event to be linked with multiple objects instead of a single case. This removes the necessity to "flatten" the event log by picking one case notion from the several potential case notions that often coexist in real-life processes.
Figure 2: Metamodel of the OCEL 2.0 format (taken from (Berti et al. 2024)).
Among the proposed implementations of OCED, the object-centric event log (OCEL) format (Ghahfarokhi et al. 2021) was designed to be more suitable for storing event logs extracted from relational databases. This model, depicted in Figure 2, is currently gaining traction in the PM community as the main contender to replace XES, supported by the availability of event logs and a wide array of software tools enabling the analysis of OCEL event logs. Table 1 shows an example log from the running example in the OCEL format. The first version of OCEL presented in (Ghahfarokhi et al. 2021) was updated in (Berti et al. 2024), which presented OCEL 2.0, solving some of the limitations of the first version of OCEL, such as the static character of object attributes, which limited the capacity of OCEL to store process context.
@@ 3.3 IoT ontologies
Recently, the focus in IoT ontologies has changed from the creation of ontologies that are as complete as possible (e.g. the Semantic Sensor Network (SSN) ontology  [^1] https://www.w3.org/TR/2017/REC-vocab-ssn-20171019/) to the development of new ontologies that are simpler and more practical (e.g. IoTStream (Elsaleh et al. 2020)). Two such ontologies are the Sensor, Observation, Sample and Actuator (SOSA) ontology (Janowicz et al. 2019) and IoTStream (Elsaleh et al. 2020).
@@ 3.3.1 SOSA
SOSA (the Sensor, Observation, Sample, and Actuator ontology) is an otology to formally define sensor networks and their observations (Janowicz et al. 2019). At its core, it is composed of three perspectives: (1) Sensors (i.e., physical or virtual) are entities capable of observing properties of features of interest. They derive measurements from other data sources. (2) Observation defines the relationship between different elements (i.e., the sensor, observed property, feature of interest, observation value). They are responsible for tracing and linking the context and origin from the environment. (3) Actuators are entities that define the state change in a feature of interest caused by executing an actuation.
@@ 3.3.2 IoTStream
IoTStream is a more specific ontology, inspired by SOSA, that focuses on the treatment of streaming data (Elsaleh et al. 2020). The main purpose of IoTStream is to use semantic technologies to bridge interoperability and heterogeneity issues. These problems arise due to the variety of devices in IoT environments, which also provide different data, making it considerably more difficult to analyze these data and gain advantages. Using IoTStream, metadata descriptions can be provided to help search for needed data by representing their properties, such as attribute values and types, finally making data analysis more accessible.
@@ 3.4 Related Work
In this section, we introduce previously developed models for IoT-enhanced event logs. In particular, Section 3.4.1 introduces the DataStream extension for XES (Mangler et al. 2023); Section 3.4.2 presents the NICE log model (Bertrand et al. 2023); and Section 3.4.3 outlines the CAIRO metamodel (Franceschetti et al. 2023c).
@@ 3.4.1 DataStream
The DataStream XES extension (Mangler et al. 2023; Ehrendorfer et al. 2023) aims at embedding different kinds of IoT sensor data into eXtensible Event Stream (XES) logs (Günther and Verbeek 2014). Therefore, it proposes to integrate additional concepts into the logs (cf. Fig. 3): stream:point (for representing single measurements), stream:multipoint (for representing multiple measurements with common attributes such as the same id or timestamp), stream:datastream (for grouping multiple points/multipoints together to link them to single events), and stream:datacontext (to provide the possibility to link sensor data to multiple events). By this, it is possible to represent IoT events in addition to process events in trace format. Furthermore, the lifecycle transition stream/data is introduced to allow recording of sensor data independently of the lifecycle events of activities and their control-flow, resulting in sensor data being available at an earlier point in time when monitoring the event stream. This procedure enables the use of the IoT sensor data for checking the conformance of process executions and allows to relate process events to corresponding IoT events and their data streams. The process and IoT events can be stored in a single file, but due to the large amount of sensor data, this is generally not the case. One advantage of DataStream is the possibility to enhance IoT-enriched event logs with semantic annotations so that process or IoT events and their origin can be semantically described in an ontology, for example. In addition, it is possible to store metadata about used data sources in the DataStream format. In contrast, there is no way to handle disambiguity in the log format. In summary, the DataStream XES extension allows for integrating IoT sensor data into XES logs and therefore enables analysis based on the log (ex-post) or based on event streams (at run-time).
Figure 3: DataStream XES Metamodel (Mangler et al. 2023)
@@ 3.4.2 NICE
NICE (Bertrand et al. 2023) aims at addressing four main challenges in IoT-enhanced event log storage: i) the granularity gap between process and IoT data, ii) the perspective convolution between control-flow and context information, iii) the scope of relevance of the IoT data and iv) the dynamicity of IoT data. To do so, NICE proposes a lightweight metamodel comprising seven main concepts, depicted in Figure 4. At its core, a NICE Event Log contains lists of Data sources, Objects and Events. An Event is related to one or several Objects. All Objects can have a collection of Properties, which represent context parameters of the process, e.g., the total amount of the order, the temperature in the fridge, the weight and the value of the parcel. Events are derived from one or several Data entries or lower-level Events which are logged by a given Data source, e.g., an information system (IS) or a sensor. IS can record IS data entries, which typically correspond to rows in a relational database, while sensors make observations, which represent individual measurements of a property of a real-world object.
Figure 4: Main concepts of the NICE model (Bertrand et al. 2023).

### v2

# Ronny Seiger5, Agnes Koschmider2, Stefanie Rinderle-Ma4, Barbara Weber5 and Estefania Serral 7

@@
An object-centric core metamodel for IoT-enhanced event logs [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
The development of Internet-of-Things (IoT) technologies has led to a widespread deployment of IoT devices by organizations to support their business processes (BPs). This trend is observed across many industries, for example, in manufacturing, where sensors and actuators pave the way for process automation (Malburg.2020_BPMAndIoT); in healthcare, where sensors continuously monitor the vital signs of patients to assist clinical decision-making and enforce best practices in patient handling (fernandez2015process); in smart spaces, where sensors and actuators support the routines of elderly people (Leotta2015ApplyingPM).
@@
As discussed by janiesch2020internet, the emergence of these so-called IoT-enhanced BPs calls for the development of dedicated business process management (BPM) techniques. These should, for example, coordinate process execution with IoT devices, respond to issues detected by sensors, and provide deeper insights related to the influence of the physical environment on process execution. To achieve these goals, new IoT-enhanced process mining (PM) techniques should be capable of analyzing IoT data together with process events.
@@
One major hurdle preventing the creation of IoT-enhanced PM techniques is the need to integrate IoT data and process event data (bertrand2022assessing; mangler2024internet). This difficulty is mainly due to the nature of IoT sensor data that is generally at a low granularity level. This in turn means that the data has to be abstracted or filtered to extract relevant information. Moreover, IoT sensor data is noisy and voluminous. This is exacerbated by the fact that IoT devices usually collect data independently of the processes. As a result, relevant IoT sensor data has to be correlated with process data to ensure a meaningful alignment between IoT observations and process events.
@@
To address this integration challenge, several models describing IoT-enhanced event log formats have recently been proposed (e.g., datastream; bertrand2023nice; franceschetti2023event). Although they share some similarities, these formats have been designed with different requirements. This is mainly due to 1) each model focusing on specific types of use cases, i.e., different types of processes and IoT devices, which generate different types of data, and 2) each model being optimized for a different data analysis task, with specific input data requirements. For example, DataStream (datastream) assumes a certain degree of process knowledge, typical of manufacturing production processes, and is intended for long-term data storage for auditing. In contrast, CAIRO (franceschetti2023event) is designed for primarily manual and more knowledge-intensive processes. Its main analysis goal is process monitoring in the absence of an information system that tracks processes. In addition to this, they are grounded on different event log standards, i.e., DataStream is an extension of the eXtensible Event Stream (XES) standard (gunther2014xes), while NICE (bertrand2023nice) follows the object-centric paradigm. This discrepancy between existing IoT-enhanced log formats hinders interoperability and data reuse in research.
@@ 2 Running example
To illustrate the need for a specialized format for IoT-enhanced event logs, this section outlines a running real-life manufacturing use case, first introduced in Bertrand2024Feb. Figure 1 shows a BPMN diagram of this process.
This use case presents the production process of a company active in the preparation of chemical products. Their production process comprises four main steps: 1) Preparing raw material and loading the tank; 2) Mixing the raw material in the tank; 3) Circulating the product through filters to remove impurities; and 4) Bottling and packing the finished product.
@@
Sensors placed around the process monitor various physical parameters affected by the process, such as the tank pressure and the flow of the product through the filters. The execution of each activity leaves a specific pattern in these IoT data, hence events and activity executions can be detected by analyzing the IoT data. For instance, a sudden peak in product flow may signal a filtering anomaly.
Early detection of flow peaks would enable process engineers to address anomalies early and to prevent the flow from fluctuating too much, which can cause degradation of the quality of the end product. However, this requires analyzing the IoT data from the flow along with process data indicating which activity is currently being executed, a task that is not straightforward. The IoT data are at a low granularity level and need to be abstracted to gain business meaning before they can be combined with process event data (mangler2024internet).
@@ 3.1 Process mining
Process mining is a method that leverages event data to analyze and enhance business processes (DataScienceInAction). It acts as a bridge between the fields of data science and process science (vanderAalst2022; DataScienceInAction). While data science typically does not prioritize processes and process science focuses mainly on process modeling rather than extracting insights from event data, process mining serves as a crucial tool that seamlessly combines the strengths of both disciplines (DataScienceInAction).
@@
    Process Discovery focuses on the automatic creation of process models derived from an event log (van2022foundations).
@@
    Conformance checking, on the other hand, aims to identify deviations between real-world process events recorded in event logs and a predefined process model (Carmona2018Conformance).
@@
    Performance analysis addresses challenges such as delays in case completion or missed deadlines. By integrating event data with process models using techniques like token-based replay and alignment methods, it uncovers inefficiencies and bottlenecks (Leemans; vanderAalst2022).
@@
    Comparative process mining relies on the analysis of multiple event logs, which may differ based on attributes such as location, time period, or resource allocation. This approach allows for the identification of critical similarities and differences between processes (vanderAalst2022; comparativepm).
@@
    In contrast to these retrospective techniques, predictive process mining combines predictive analytics with process mining methods to forecast future events, outcomes, and remaining time in support of proactive decision-making (Schulte2017Mar; vanderAalst2022; DiFrancescomarino2022Jun).
@@
    Finally, action-oriented process mining bridges the gap between analysis and implementation by translating insights into actionable steps to enhance operational efficiency, e.g. using robotic process automation (Park2022Jul; vanderAalst2022).
@@ 3.1.1 Process Mining in IoT
IoT devices generate vast amounts of data reflecting the physical context in which processes operate, offering a rich resource for process mining. Using continuous streams of data from interconnected devices, process mining can move beyond traditional applications to provide context-sensitive, real-time insights into processes, opening new possibilities for advanced techniques (Leotta2015ApplyingPM). For instance, process mining is increasingly employed in IoT-based systems to detect misbehavior and potential attacks (Hemmer), as well as to diagnose discrepancies between the designed IoT system and its actual usage. This facilitates anomaly detection and ensures that systems operate as intended (Yamaguchi).
The time series data produced by IoT devices often includes detailed, low-level information about processes and their executions. This enables the discovery of processes at various levels of abstraction, enhancing event log-driven process discovery by embedding IoT concepts directly into process models (Elkodssi2023Mar). Moreover, IoT data provides valuable contextual information about the physical environments in which processes occur. Incorporating this context into process mining allows for a more in-depth understanding of the factors influencing process execution, leading to more informed decision-making (Bertrand2024Feb).
However, integrating IoT data into process mining introduces significant challenges. First, there is a substantial gap between the granular data in sensor logs and the higher-level actions needed for process mining, further complicated by the uncertainty and noise inherent in IoT data, which may lead to ambiguity (franceschetti2023characterisation). Second, modeling human behavior is complex due to its unstructured and flexible nature, requiring adaptations such as fuzzy mining or declarative approaches to account for variability and probabilistic patterns. Finally, automatic log segmentation is particularly challenging as sensor logs lack predefined traces and must account for varied routines, temporal constraints, and multi-user interactions, necessitating innovative methods for accurately defining and differentiating human habits (Leotta2015ApplyingPM).
@@ 3.2.1 XES
The current standard event log model is the eXtensible Event Stream (XES) format, an XML-based model that mainly consists of the notions of event, case, and log (gunther2014xes). It proposes standard attribute types to contextualize the events, e.g. the resource executing an activity, the cost of an activity, etc. A standard activity lifecycle is defined together with XES, based on which the status of an activity can be mapped with events relating to this activity. XES also allows the definition of new data attribute types through extensions.
@@ 3.2.2 OCEL/OCED
Recently, the gain in maturity of the PM field has increased the urge to create alternative models to the original XES standard. Multiple propositions that relax some assumptions of XES and allow for more flexibility in event data storage have been presented, e.g., in popova2015artifact; ghahfarokhi2021ocel; berti2024ocel; fahland2024towards. In this context, the object-centric event data (OCED) initiative (fahland2024towards) led by the IEEE task force in process mining (TFPM) aims at updating the XES format used by the community with a more flexible standard for event data. The main innovation of OCED is the introduction of the concept of object, which generalizes the notion of case by allowing one event to be linked with multiple objects instead of a single case. This removes the necessity to "flatten" the event log by picking one case notion from the several potential case notions that often coexist in real-life processes.
Figure 2: Metamodel of the OCEL 2.0 format (taken from (berti2024ocel)).
Among the proposed implementations of OCED, the object-centric event log (OCEL) format (ghahfarokhi2021ocel) was designed to be more suitable for storing event logs extracted from relational databases. This model, depicted in Figure 2, is currently gaining traction in the PM community as the main contender to replace XES, supported by the availability of event logs and a wide array of software tools enabling the analysis of OCEL event logs. Table 1 shows an example log from the running example in the OCEL format. The first version of OCEL presented in (ghahfarokhi2021ocel) was updated in (berti2024ocel), which presented OCEL 2.0, solving some of the limitations of the first version of OCEL, such as the static character of object attributes, which limited the capacity of OCEL to store process context.
@@ 3.3 IoT ontologies
Recently, the focus in IoT ontologies has changed from the creation of ontologies that are as complete as possible (e.g. the Semantic Sensor Network (SSN) ontology  [^2] https://www.w3.org/TR/2017/REC-vocab-ssn-20171019/) to the development of new ontologies that are simpler and more practical (e.g. IoTStream (elsaleh2020iot)). Two such ontologies are the Sensor, Observation, Sample and Actuator (SOSA) ontology (janowicz2019sosa) and IoTStream (elsaleh2020iot).
@@ 3.3.1 SOSA
SOSA (the Sensor, Observation, Sample, and Actuator ontology) is an otology to formally define sensor networks and their observations (janowicz2019sosa). At its core, it is composed of three perspectives: (1) Sensors (i.e., physical or virtual) are entities capable of observing properties of features of interest. They derive measurements from other data sources. (2) Observation defines the relationship between different elements (i.e., the sensor, observed property, feature of interest, observation value). They are responsible for tracing and linking the context and origin from the environment. (3) Actuators are entities that define the state change in a feature of interest caused by executing an actuation.
@@ 3.3.2 IoTStream
IoTStream is a more specific ontology, inspired by SOSA, that focuses on the treatment of streaming data (elsaleh2020iot). The main purpose of IoTStream is to use semantic technologies to bridge interoperability and heterogeneity issues. These problems arise due to the variety of devices in IoT environments, which also provide different data, making it considerably more difficult to analyze these data and gain advantages. Using IoTStream, metadata descriptions can be provided to help search for needed data by representing their properties, such as attribute values and types, finally making data analysis more accessible.
@@ 3.4 Related Work
In this section, we introduce previously developed models for IoT-enhanced event logs. In particular, Section 3.4.1 introduces the DataStream extension for XES (datastream); Section 3.4.2 presents the NICE log model (bertrand2023nice); and Section 3.4.3 outlines the CAIRO metamodel (franceschetti2023event).
@@ 3.4.1 DataStream
The DataStream XES extension (datastream; Ehrendorfer.2023_IoPTRepository) aims at embedding different kinds of IoT sensor data into eXtensible Event Stream (XES) logs (gunther2014xes). Therefore, it proposes to integrate additional concepts into the logs (cf. Fig. 3): stream:point (for representing single measurements), stream:multipoint (for representing multiple measurements with common attributes such as the same id or timestamp), stream:datastream (for grouping multiple points/multipoints together to link them to single events), and stream:datacontext (to provide the possibility to link sensor data to multiple events). By this, it is possible to represent IoT events in addition to process events in trace format. Furthermore, the lifecycle transition stream/data is introduced to allow recording of sensor data independently of the lifecycle events of activities and their control-flow, resulting in sensor data being available at an earlier point in time when monitoring the event stream. This procedure enables the use of the IoT sensor data for checking the conformance of process executions and allows to relate process events to corresponding IoT events and their data streams. The process and IoT events can be stored in a single file, but due to the large amount of sensor data, this is generally not the case. One advantage of DataStream is the possibility to enhance IoT-enriched event logs with semantic annotations so that process or IoT events and their origin can be semantically described in an ontology, for example. In addition, it is possible to store metadata about used data sources in the DataStream format. In contrast, there is no way to handle disambiguity in the log format. In summary, the DataStream XES extension allows for integrating IoT sensor data into XES logs and therefore enables analysis based on the log (ex-post) or based on event streams (at run-time).
Figure 3: DataStream XES Metamodel (datastream)
@@ 3.4.2 NICE
NICE (bertrand2023nice) aims at addressing four main challenges in IoT-enhanced event log storage: i) the granularity gap between process and IoT data, ii) the perspective convolution between control-flow and context information, iii) the scope of relevance of the IoT data and iv) the dynamicity of IoT data. To do so, NICE proposes a lightweight metamodel comprising seven main concepts, depicted in Figure 4. At its core, a NICE Event Log contains lists of Data sources, Objects and Events. An Event is related to one or several Objects. All Objects can have a collection of Properties, which represent context parameters of the process, e.g., the total amount of the order, the temperature in the fridge, the weight and the value of the parcel. Events are derived from one or several Data entries or lower-level Events which are logged by a given Data source, e.g., an information system (IS) or a sensor. IS can record IS data entries, which typically correspond to rows in a relational database, while sensors make observations, which represent individual measurements of a property of a real-world object.
Figure 4: Main concepts of the NICE model (bertrand2023nice).